OpenStudy (jotopia34):

Integral question: Please!! the Integral of 1/sqrt(2x-x^2) dx

OpenStudy (jotopia34):

\[\int\limits1/({2x-x^2)}\]

pooja195 (pooja195):

∫ 1 / sqrt(2x-x^2) dx Let's complete the square. ∫ 1 / sqrt(-(x^2-2x+1) + 1) dx ∫ 1 / sqrt(1 - (x-1)^2) dx Let, u = x-1 du = dx ∫ 1 / sqrt(1 - u^2) du arcsin(u) + C arcsin(x-1) + C

OpenStudy (jotopia34):

pellet, i did sorry

OpenStudy (jotopia34):

okay, complete the square it is. Only I don't remember how...:( please help pooja

pooja195 (pooja195):

This is the original problem. 4x2 – 2x – 5 = 0 Move the loose number over to the other side. 4x2 – 2x = 5 Divide through by whatever is multiplied on the squared term. Take half of the coefficient (don't forget the sign!) of the x-term, and square it. Add this square to both sides of the equation. Convert the left-hand side to squared form, and simplify the right-hand side. (This is where you use that sign that you kept track of earlier. You plug it into the middle of the parenthetical part.) Square-root both sides, remembering the "±" on the right-hand side. Simplify as necessary. Solve for "x =". Remember that the "±" means that you have two values for x.

pooja195 (pooja195):

ps just an example

OpenStudy (jotopia34):

Ok, i see what you are saying, just had to refresh the memory. Thx again